# Advanced Encryption Standard AI artificial intelligence

### v2

@@
Simon Klix $0$-0002-9369-2901 MPI-SPBochumGermany Nils Albartus $0$-0003-2449-1134 MPI-SPBochumGermany Julian Speith $0$-0002-8408-8518 MPI-SPBochumGermany Paul Staat $0$-0002-7539-4847 MPI-SPBochumGermany Jörn Langheinrich $0$-0002-8583-5503 MPI-SPBochumGermany Alice Verstege 0009-0005-2807-0806 Ruhr University BochumBochumGermany Annika Wilde $0$-0002-6092-7866 Ruhr University BochumBochumGermany Daniel Lammers $0$-0002-9134-8568 Ruhr University BochumBochumGermany Christian Kison $0$-0002-5830-7692 BundeskriminalamtWiesbadenGermany Sebastian Sester-Wehle $0$-0002-4938-2084 BundeskriminalamtWiesbadenGermany Daniel Holcomb $0$-0002-2052-9820 UMass AmherstAmherstMAUSA  and  Christof Paar $0$-0001-8681-2277 MPI-SPBochumGermany
@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ACM Conference on Computer and Communications Security; Oct. 14-18, 2024; Salt Lake City, UT, USAisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Hardware reverse engineering
@@ 1. Introduction
Modern nanometer-scale ICs come with a natural defense against reverse engineering due to the high cost and labor-intensive nature of the netlist extraction process (Lippmann et al., 2019; Fyrbiak et al., 2017; Quadir et al., 2016; Torrance and James, 2011, 2009)—a financial barrier that deters all but the most resourceful entities, such as nation-states. However, the bar for FPGA reverse engineering is much lower. To this end, an attacker would need to access the bitstream and map it to a gate-level netlist. This process is already well understood despite the proprietary nature of the bitstream format (Ender et al., 2019; Ziener et al., 2006; Note and Rannaud, 2008; Benz et al., 2012; Ding et al., 2013; Pham et al., 2017; Note, 2008; Alliance, 2023b). FPGA vendors commonly provide cryptographic bitstream protections as a safeguard for valuable IP. Nonetheless, these measures have repeatedly been demonstrated to be susceptible to side-channel attacks (Moradi et al., 2011, 2012; Moradi and Schneider, 2016; Swierczynski et al., 2015b; Tajik et al., 2017) or protocol flaws (Ender et al., 2020, 2022). Many low-cost FPGAs do not even offer such protections. Even if these protective measures are available and resist such attacks, they must be proactively enabled by the FPGA developer. Hence, we assume access to the plaintext bitstream, and thus the gate-level netlist, is generally achievable in practice.
@@ 1. RQ1:
To answer RQ1, we demonstrate the steps and efforts required to extract IP from an FPGA implementation through a case study on a Lattice iCE40 FPGA found within the iPhone 7. Apple refers to this FPGA as Maggie in the iPhone firmware. To this end, we combine existing techniques and propose new approaches to recover an algorithmic description of the target FPGA implementation, effectively extracting the IP controlling the iPhone’s Taptic Engine, see Section 3. In addition to evaluating existing techniques and proposing novel approaches to FPGA reverse engineering, we provide unique insights that foster developing better defenses against IP theft. Based on our findings, a second research question arises:
@@ 2.1. Attacker Model
As is typical in reverse engineering, the attacker may have already gathered public information on the target device, such as documentation or patents. This leaves them with knowledge on the system’s connectivity and interaction with the FPGA, along with a general understanding of its purpose, but lacking detailed knowledge of the implementation and the intricacies of the algorithm. Ultimately, the attacker seeks to reconstruct a detailed algorithmic representation that allows him to extract, understand, and modify the IP contained within it. Thus, our attacker model explicitly goes beyond simply copying an FPGA bitstream or netlist.
@@ 3. C3:
    Synthesizer Optimizations. Synthesizer optimizations can result in similar operations on HDL level looking vastly different in a gate-level netlist. Logic may be merged across module boundaries and gates or even entire sub-circuits may be duplicated to facilitate a high fan-out. Such optimizations always depend on the available standard cells, timing and area constraints, routing requirements, and varying optimization strategies.
@@ 3. Case Study on iPhone 7
From analyzing leaked PCB schematics, the iPhone firmware, and patents granted to Apple (Hajati, 2021; Hajati and Patel, 2018), we concluded that Maggie controls the Taptic Engine, and determined its I/O connectivity and communication interfaces, see Figure 1. This alone does not enable effective IP theft, as we still lack implementation details on the instantiated signal processing algorithms and their parameters. Assuming the role of the attacker, we therefore need to investigate the circuit implemented on Maggie to fully recover Apple’s Taptic Engine IP. To this end, we split the iPhone FPGA reverse engineering process into four steps, see Figure 2.
@@ 3.1. Step 1: Netlist Recovery
SRAM-based FPGAs such as Lattice iCE40 devices store the bitstream externally and are configured during startup. Hence, an attacker can read the bitstream from external memory, intercept it during transmission, or recover it from device firmware. The Maggie bitstream is shipped with the openly available iPhone firmware, hence we extracted it from the file system. The original bitstream was published in September 2016 and was updated with iOS 10.2 in December 2016. As it remains unchanged ever since, our analysis deals with this latest version.
@@ 3.2. Step 2: Word-Level Reconstruction
MUXes in particular provide structure to the data path. On FPGAs, however, their functionality is implemented as part of more complex Boolean functions within LUTs. Hence, decomposing complex LUTs into small building blocks such as MUXes can aid manual and automated reverse engineering, e.g., by supporting structural analysis. To recover nested MUXes and better understand the data path (see Figure 5), we analyzed each LUT’s Boolean function and searched for inputs that behave like a select. When discovering a MUX in a LUT, we replaced the LUT with a MUX and reconstruct the surrounding logic from primitive gates. In total, we replaced 549 LUTs with primitive gates, removed 594 buffers, extracted 1619 MUXes and 1161 primitive gates, removed 353 duplicate gates, and removed 51 additional gates using other simplifications.
@@
To recover word-level arithmetic operations, we first located carry chains in the netlist by scanning for connected carry gates. We then assembled arithmetic structure candidates by adding surrounding logic to the carry chains. Since a structure may be influenced by control inputs, we constructed structural candidate variants to account for different control behaviors. Operand membership and the order of input and output signals were derived from the structure and connectivity of a candidate’s gates. For example, the order of inputs was reconstructed by identifying the carry gate of the chain that a signal is connected to. Finally, we checked each candidate against predefined satisfiability modulo theories (SMT) models, e.g., $A\pm B$ for adders or $A\pm n$ for counters incrementing by a constant $n$.
@@ 3.4. Step 4: High-Level Sense-Making
Based on the recovered algorithm, we proceeded with a semantic analysis that is essential for meaningful interaction with Apple’s IP. Generating a Python script implementing the algorithm proved crucial, since comprehending the algorithm required domain knowledge that was beyond our FPGA reverse engineering team. This way, we could share the Python script containing the reconstructed algorithm with a signal processing expert without requiring them to have any knowledge of the reverse engineering process.
@@
Maggie is a central component of the iPhone 7 haptics subsystem to which Apple refers as Taptic Engine. We found that it controls the excitation of the  linear resonant actuator (LRA) to produce a vibration which is perceived by the user as a tactile stimulus. For instance, the Taptic Engine creates the sensation of a physical button push for the non-mechanical home button of the iPhone 7. To provide such refined tactile feedback, a closed-loop control system is utilized which makes the LRA accurately follow a desired movement. Here, Maggie implements the closed-loop motion controller, see Figure 9 for a simplified block diagram, which conditions the drive signals applied to the actuator (via the motor driver) based on sensor feedback from the actuator itself.
@@
To generate the LRA control signals, Maggie operates on two feedback signals from the LRA, a reference signal previously loaded into BRAM, and values of intermediate signals at previous sampling instants. From the algorithmic representation, we identified several classical biquad filter stages implemented in a direct form I structure (Oppenheim and Schafer, 2014). Upon closer inspection, we found many typical DSP building blocks, e.g., sub-sampling and sample-and-hold blocks for sample rate conversion, bit shifts and truncations for fixed-point arithmetic, saturation stages to prevent overflows, and filter implementations. Furthermore, we recognized even subtle details such as the use of fraction saving (Yates and Lyons, 2008). Putting it all together, we found three main processing stages: $(i)$ input signal conditioning and application of calibration, $(ii)$ a state observer to track the actuator dynamics, and $(iii)$ the closed-loop controller to minimize errors between the actual and the desired LRA movement. For details on the recovered DSP operations, see Appendix B.
@@ 4.1. Netlist Pre-Processing
So far, our pre-processing steps were limited to the 4-input LUTs found on Lattice iCE40 FPGAs. While some techniques such as the detection of duplicate gates could easily be extended to cover arbitrary combinational gates, this is not true for our LUT decomposition. Here, our approach works reasonably well for 4-input LUTs but does not scale to six inputs. Hence, for generalization, we re-synthesize the combinational subgraphs between sequential gates using open-source EDA tools instead of analyzing each LUT’s Boolean function. To this end, we constrain the synthesizer to a custom gate library comprising only the desired combinational gates such as AND, OR, XOR, and MUX of various sizes. Compared to decomposing the combinational parts of the netlist into a pure and-or-invert (AOI) graph, this approach allows us to delegate multiplexer (MUX) detection to the synthesizer by incentivizing the selection of MUX gates over alternatives by artificially reducing their size and optimizing for area. This way, we can even identify MUXes within complex LUT configurations. By nature, this re-synthesis approach is independent of the underlying FPGA architecture and target implementation.
@@ 4.2. Arithmetic Structures
We verified functionality under at least one control assignment for 94% of the carry chains across our benchmarks, see Table 2. Issues only arose when carry chains where strongly interwoven with surrounding control and data-path logic, which sometimes prevented extraction of correct structural candidates. Depending on the benchmark, we automatically verified that 1% to 91% (average: 32%) of all combinational gates implement an arithmetic operation. The variation in coverage is the result of different kinds of benchmark implementations. For signal processing designs like the canny_edge_detector and the fft64, the data path mostly comprises of arithmetic operations. CPU implementations such as ibex, icicle, and simple_risc_v contain fewer carry chains and more control logic, which explains lower numbers in detected arithmetic operations and overall combinational coverage.
@@ 4.3. Data Path
The idea behind the dataflow analysis methodology of DANA (Albartus et al., 2020) is, by nature, independent of the target gate type. In our case study, we leveraged DANA to identify registers and developed a separate, but related, technique to group MUX gates into word-level structures. Similar to DANA, we used control signals as well as predecessor and successor gates to identify word-level MUXes. Accordingly, to generalize our data-path analysis, we extended DANA to take on arbitrary, user-defined gate types, including MUXes, instead of refining our initial approach. In addition, we added the option to provide known word-level structures of any kind to DANA to support word-level recovery of the target gate type. For example, in our case study DANA could only search for registers and only previously identified registers could be provided as known groups. Now, DANA can also reconstruct word-level MUXes while taking previously identified registers and arithmetic operations into account.
@@
To evaluate the success of register recovery, we followed the evaluation proposed for DANA (Albartus et al., 2020). We generated a ground truth by leveraging symbols left in the gate-level netlists of our benchmark by the synthesizer. Then, we use normalized mutual information (NMI) and purity as metrics to compare our refined approach against the original DANA implementation. To this end, we fed known word-level structures such as BRAMs, DSPs, and arithmetic operations to DANA as additional predefined knowledge to take into account during analysis. This marginally improved NMI and purity for most benchmarks, raising the NMI by up to 0.14 the the purity by up to 0.30, see Table 2.
We also employed DANA to recover word-level MUXes. However, by nature, no ground truth for MUXes exists, as they are implicitly created from HDL control flow structures by the synthesizer and then implemented in LUTs alongside other logic. Therefore, it is, at best, ambiguous which HDL constructs would be translated into a MUX and accurately measuring the quality of our results is impossible. Still, we provide an intuition for their quality by investigating the sizes of recovered MUXes in Appendix D.

### v3

@@
Simon Klix $0$-0002-9369-2901 MPI-SPBochumGermany Nils Albartus $0$-0003-2449-1134 MPI-SPBochumGermany Julian Speith $0$-0002-8408-8518 MPI-SPBochumGermany Paul Staat $0$-0002-7539-4847 MPI-SPBochumGermany Alice Verstege 0009-0005-2807-0806 MPI-SPBochumGermany Annika Wilde $0$-0002-6092-7866 Ruhr University BochumBochumGermany Daniel Lammers $0$-0002-9134-8568 Ruhr University BochumBochumGermany Jörn Langheinrich $0$-0002-8583-5503 MPI-SPBochumGermany Christian Kison $0$-0002-5830-7692 BundeskriminalamtWiesbadenGermany Sebastian Sester-Wehle $0$-0002-4938-2084 BundeskriminalamtWiesbadenGermany Daniel Holcomb $0$-0002-2052-9820 UMass AmherstAmherstMAUSA  and  Christof Paar $0$-0001-8681-2277 MPI-SPBochumGermany
@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security; October 14–18, 2024; Salt Lake City, UT, USAbooktitle: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security (CCS ’24), October 14–18, 2024, Salt Lake City, UT, USAdoi: 10.1145/3658644.3690235isbn: 979-8-4007-0636-3/24/10ccs: Security and privacy Hardware reverse engineering
@@ 1. Introduction
Modern nanometer-scale ICs come with a natural defense against reverse engineering due to the high cost and labor-intensive nature of the netlist extraction process (Lippmann et al., 2019; Fyrbiak et al., 2017; Quadir et al., 2016; Torrance and James, 2011, 2009)—a financial barrier that deters all but the most resourceful entities, such as nation-states. However, the bar for FPGA reverse engineering is much lower. To this end, an attacker would need to access the bitstream and map it to a gate-level netlist. This process is already well understood despite the proprietary nature of the bitstream format (Ender et al., 2019; Ziener et al., 2006; Note and Rannaud, 2008; Benz et al., 2012; Ding et al., 2013; Pham et al., 2017; Note, 2008; Alliance, 2023b). FPGA vendors commonly provide cryptographic bitstream protections as a safeguard for valuable IP. Nonetheless, these measures have repeatedly been demonstrated to be susceptible to side-channel attacks (Moradi et al., 2011a, 2012; Moradi and Schneider, 2016; Swierczynski et al., 2015b; Tajik et al., 2017) or protocol flaws (Ender et al., 2020, 2022). Many low-cost FPGAs do not even offer such protections. Even if these protective measures are available and resist such attacks, they must be proactively enabled by the FPGA developer. Hence, we assume access to the plaintext bitstream, and thus the gate-level netlist, is generally achievable in practice.
@@ 1. RQ1:
To answer RQ1, we demonstrate the steps and efforts required to extract IP from an FPGA implementation through a case study on a Lattice iCE40 FPGA found within the iPhone 7. To this end, we combine existing techniques and propose new approaches to recover an algorithmic description of the target FPGA implementation, effectively extracting the IP controlling the iPhone’s Taptic Engine, see Section 3. In addition to evaluating existing techniques and proposing novel approaches to FPGA reverse engineering, we provide unique insights that foster developing better defenses against IP theft. While the iPhone 7 was released in 2016, Lattice iCE40 FPGAs are still used in many other devices such as Apple Vision Pro, Apple TV 4K, HTC Vive Pro, Samsung Galaxy S5, and Pebble Time. As the bitstreams of all Lattice iCE40 FPGAs follow the same structure, our tooling could be directly applied to the bitstreams of these other devices to obtain their gate-level netlists. Subsequently, the netlist reverse engineering techniques presented in this paper could be used to analyze the extracted netlist right away.
Please note that Apple refers to the FPGA inside iPhone 7 as Maggie in the firmware, and we will also use this name throughout our paper. Based on our findings, a second research question arises:
@@ 2.1. Attacker Model
As is typical in reverse engineering, the attacker may have already gathered public information on the target device, such as documentation or patents. This leaves them with knowledge of the system’s connectivity and interaction with the FPGA, along with a general understanding of its purpose, but lacking detailed knowledge of the implementation and the intricacies of the algorithm. Ultimately, the attacker seeks to reconstruct a detailed algorithmic representation that allows him to extract, understand, and modify the IP contained within it. Thus, our attacker model explicitly goes beyond simply copying an FPGA bitstream or netlist.
@@ 3. C3:
    Synthesizer Optimizations. Synthesizer optimizations can result in similar operations on HDL level looking vastly different in a gate-level netlist. Logic may be merged across module boundaries and gates, or entire sub-circuits may be duplicated to facilitate a high fan-out. Such optimizations always depend on the available standard cells, timing and area constraints, routing requirements, and varying optimization strategies.
@@ 3. Case Study on iPhone 7
From analyzing leaked PCB schematics, the iPhone firmware, and patents granted to Apple (Hajati, 2021; Hajati and Patel, 2018), we concluded that Maggie controls the Taptic Engine, and determined its I/O connectivity and communication interfaces, see Figure 1. We further discovered that the FPGA itself is programmed on start-up and is not used as a re-programmable accelerator; it always implements the same control logic of the Taptic Engine, and the design does not differ between applications. The FPGA is controlled by the A10 system-on-chip (SoC) and triggered by the OS to generate haptic feedback, e.g., to simulate home button presses. However, this knowledge alone does not enable effective IP theft, as we still lack implementation details on the instantiated signal processing algorithms and their parameters. Assuming the role of the attacker, we therefore need to investigate the circuit implemented on Maggie to fully recover Apple’s Taptic Engine IP. To this end, we split the iPhone FPGA reverse engineering process into four steps, see Figure 2.
@@ 3.1. Step 1: Netlist Recovery
SRAM-based FPGAs such as Lattice iCE40 devices store the bitstream externally and are configured during startup. Hence, an attacker can read the bitstream from external memory, intercept it during transmission, or recover it from device firmware. The Maggie bitstream is shipped with the openly available iPhone firmware, hence we extracted it from the file system. The original bitstream was published in September 2016 and was updated with iOS 10.2 in December 2016. As it has remained unchanged ever since, our analysis deals with this latest version. As Lattice iCE40 FPGAs do not offer any bitstream protections such as encryption, we were directly faced with then plaintext bitstream. This lack of security features can occasionally be observed in low-cost FPGAs, leaving the IP they implement mostly unprotected. Please refer to Appendix F for a list of FPGAs that do (not) support bitstream encryption.
@@ 3.2. Step 2: Word-Level Reconstruction
MUXes in particular provide structure to the data path. On FPGAs, however, their functionality is implemented as part of more complex Boolean functions within LUTs. Hence, decomposing complex LUTs into small building blocks such as MUXes can aid manual and automated reverse engineering, e.g., by supporting structural analysis. To recover nested MUXes and better understand the data path (see Figure 5), we analyzed each LUT’s Boolean function and searched for inputs that behave like a select. When discovering a MUX in a LUT, we replaced the LUT with a MUX and reconstructed the surrounding logic from primitive gates. In total, we replaced 549 LUTs with primitive gates, removed 594 buffers, extracted 1619 MUXes and 1161 primitive gates, removed 353 duplicate gates, and removed 51 additional gates using other simplifications.
@@
To recover word-level arithmetic operations, we first located carry chains in the netlist by scanning for connected carry gates. We then assembled arithmetic structure candidates by adding surrounding logic to the carry chains. Since a structure may be influenced by control inputs, we constructed structural candidate variants to account for different control behaviors. Operand membership and the order of input and output signals were derived from the structure and connectivity of a candidate’s gates. For example, the order of inputs was reconstructed by identifying the carry gate of the chain to which a signal is connected. Finally, we checked each candidate against predefined satisfiability modulo theories (SMT) models, e.g., $A\pm B$ for adders or $A\pm n$ for counters incrementing by a constant $n$.
@@ 3.4. Step 4: High-Level Sense-Making
Based on the recovered algorithm, we proceeded with a semantic analysis that is essential for meaningful interaction with Apple’s IP. Generating a Python script implementing the algorithm proved crucial since comprehending the algorithm required domain knowledge that was beyond our FPGA reverse engineering team. This way, we could share the Python script containing the reconstructed algorithm with a signal processing expert without requiring them to have any knowledge of the reverse engineering process.
@@
Maggie is a central component of the iPhone 7 haptics subsystem to which Apple refers as Taptic Engine. We found that it controls the excitation of the  linear resonant actuator (LRA) to produce a vibration which the user perceives as a tactile stimulus. For instance, the Taptic Engine creates the sensation of a physical button push for the non-mechanical home button of the iPhone 7. To provide such refined tactile feedback, a closed-loop control system is utilized, which makes the LRA accurately follow a desired movement. Here, Maggie implements the closed-loop motion controller, see Figure 9 for a simplified block diagram, which conditions the drive signals applied to the actuator (via the motor driver) based on sensor feedback from the actuator itself.
@@
To generate the LRA control signals, Maggie operates on two feedback signals from the LRA, a reference signal previously loaded into BRAM, and values of intermediate signals at previous sampling instants. From the algorithmic representation, we identified several classical biquad filter stages implemented in a direct form I structure (Oppenheim and Schafer, 2014). Upon closer inspection, we found many typical DSP building blocks, e.g., sub-sampling and sample-and-hold blocks for sample rate conversion, bit shifts and truncations for fixed-point arithmetic, saturation stages to prevent overflows, and filter implementations. Furthermore, we recognized subtle details such as the use of fraction saving (Yates and Lyons, 2008). We could even identify the individual building blocks along with their value-exact parametrization, including filter and control-loop coefficients and initialization values—fully characterizing the entire signal processing chain. Putting it all together, we found three main processing stages: $(i)$ input signal conditioning and application of calibration, $(ii)$ a state observer to track the actuator dynamics, and $(iii)$ the closed-loop controller to minimize errors between the actual and the desired LRA movement. For additional details on the recovered DSP operations, see Appendix B.
@@ 4.1. Netlist Pre-Processing
So far, our pre-processing steps were limited to the 4-input LUTs found on Lattice iCE40 FPGAs. While some techniques such as the detection of duplicate gates could easily be extended to cover arbitrary combinational gates, this is not true for our LUT decomposition. Here, our approach works reasonably well for 4-input LUTs but does not scale to six inputs. Hence, for generalization, we re-synthesize the combinational subgraphs between sequential gates using the open-source EDA tool Yosys (Shah et al., 2019) instead of analyzing each LUT’s Boolean function. To this end, we constrain the synthesizer to a custom gate library comprising only the desired combinational gates such as INV, BUF, AND, OR, XOR, XNOR, and MUX of various sizes. Compared to decomposing the combinational parts of the netlist into a pure and-or-invert (AOI) graph, this approach allows us to delegate multiplexer (MUX) detection to the synthesizer by incentivizing the selection of MUX gates over alternatives by artificially reducing their size and optimizing for area. This way, we can even identify MUXes within complex LUT configurations. Another benefit of using only primitive gates is that it improves the efficiency of structural analysis. Allowing more complex combinational gates during re-synthesis can result in the synthesizer merging interconnected logic into fewer gates, which obscures module boundaries and hampers structural analysis. By nature, this re-synthesis approach is independent of the underlying FPGA architecture and target implementation.
@@ 4.2. Arithmetic Structures
We verified functionality under at least one control assignment for 94% of the carry chains across our benchmarks, see Table 2. Issues only arose when carry chains were strongly interwoven with surrounding control and data-path logic, which sometimes prevented the extraction of correct structural candidates. Depending on the benchmark, we automatically verified that 1% to 91% (average: 32%) of all combinational gates implement an arithmetic operation. The variation in coverage is the result of different kinds of benchmark implementations. For signal processing designs like the canny_edge_detector and the fft64, the data path mostly comprises of arithmetic operations. CPU implementations such as ibex, icicle, and simple_risc_v contain fewer carry chains and more control logic, which explains lower numbers in detected arithmetic operations and overall combinational coverage.
@@ 4.3. Data Path
The idea behind the dataflow analysis methodology of DANA (Albartus et al., 2020) is, by nature, independent of the target gate type. In our case study, we leveraged DANA to identify registers and developed a separate but related technique to group MUX gates into word-level structures. Similar to DANA, we used control signals as well as predecessor and successor gates to identify word-level MUXes. Accordingly, to generalize our data-path analysis, we extended DANA to take on arbitrary, user-defined gate types, including MUXes, instead of refining our initial approach. In addition, we added the option to provide known word-level structures of any kind to DANA to support word-level recovery of the target gate type. For example, in our case study DANA could only search for registers, and only previously identified registers could be provided as known groups. Now, DANA can also reconstruct word-level MUXes while taking previously identified registers and arithmetic operations into account.
@@
To evaluate the success of register recovery, we followed the evaluation proposed for DANA (Albartus et al., 2020). We generated a ground truth by leveraging symbols left in the gate-level netlists of our benchmark by the synthesizer. Then, we use normalized mutual information (NMI) and purity as metrics to compare our refined approach against the original DANA implementation. To this end, we fed known word-level structures such as BRAMs, DSPs, and arithmetic operations to DANA as additional predefined knowledge to take into account during analysis. This marginally improved NMI and purity for most benchmarks, raising the NMI by up to 0.14 and the purity by up to 0.30, see Table 2.
We also employed DANA to recover word-level MUXes. However, by nature, no ground truth for MUXes exists, as they are implicitly created from HDL control flow structures by the synthesizer and then implemented in LUTs alongside other logic. Therefore, it is, at best, ambiguous which HDL constructs would be translated into a MUX, and accurately measuring the quality of our results is impossible. Still, we provide an intuition for their quality by investigating the sizes of recovered MUXes in Appendix D.